Hello, i was wondering if i could be walked through this and some other problems, it is: 3+2ln x = 10
Answer by lwsshak3(11628)

3+2ln x = 10
2ln x = 7
ln x=7/2
x=e^7/2 (base(e) raised to log of number(7/2)=number(x)(exponential form)